Suppose u, v V and ||u|| = ||v|| = 1 with < u, v > = 1.

Prove that u = v.

Please assist.


1
Expert's answer
2021-06-21T09:48:50-0400

Any two vectors "u" and "v" will be equal if they are collinear, codirected, and have the same magnitude.


Given "||u||=||v||=1, (u, v)=1."

Then


"(u,v)=||u||\cdot||v||\cdot \cos\widehat{(u,v)}"

"1=1\cdot1\cdot\cos \widehat{(u,v)}=>\cos \widehat{(u,v)}=1"

"=>\widehat{(u,v)}=0"

If the angle between two non-zero vectors equal zero , then the vectors are parallel in the same direction.

We have two non-zero, collinear, codirected, unit vectors "u" and "v."

Therefore the vectors "u" and "v" are equal


"u=v"